OpenStudy (anonymous):

if P(x) = x+ 2, what is P(a+1)?

jigglypuff314 (jigglypuff314):

plug in (a+1) for x so P(a+1) = (a+1)+2 = a + 1 + 2 then simplify and that's your answer :)

OpenStudy (anonymous):

thank you thank you :)

jigglypuff314 (jigglypuff314):

you're welcome :)